OpenStudy (anonymous):

plse help me solve this sum sum no. 16

terenzreignz (terenzreignz):

Okay, let x be the time (in hours) that it walked and y be the time (in hours) that it trotted. Then you have 9x + 15y = 11

terenzreignz (terenzreignz):

Where x and y are both between 0 and 1. This is about as far as I can go...

terenzreignz (terenzreignz):

oh yeah, sorry, x + y = 1 and you have a system :D

terenzreignz (terenzreignz):

9x + 15y = 11 x + y = 1 And just solve for x and y :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

y=1/3 x=2/3

terenzreignz (terenzreignz):

no that's not it :D

terenzreignz (terenzreignz):

Let's work with the second equation first.... x + y = 1 Now, let's subtract x from both sides to yield y = 1 - x You now have an expression of y in terms of x. So what you do is replace the y in the first equation 9x + 15y = 11 with 1 - x and then solve for x.

OpenStudy (anonymous):

yes, x= 2/3 and y = 1/3 is coming

terenzreignz (terenzreignz):

oh sorry, I misread your first post of your answer as x = 1/3 and y = 2/3 Good job... sorry for the mix-up T.T

OpenStudy (anonymous):

@amistre64 i need ur help

terenzreignz (terenzreignz):

huh? you've already solved it, haven't you?

OpenStudy (amistre64):

i cant see the details of the question. the link is a book cover

terenzreignz (terenzreignz):

Second response was a question though :D

OpenStudy (amistre64):

9x + 15y = 11, find the gcd(9,15)

OpenStudy (amistre64):

i believe its 3 9x + 15y = 3 can be solved. 15 = 9(1) + 6 9 = 6(1) + 3 <-- gcd = 3 6 = 3(2) ... 0 0 1 0-1(1) = -1 1-1(-1) = 2 ; x= 2 9(2) + 15(-1) = 3 and y = -1 multiply it all by 11/3 9(22/3) + 15(-11/3) = 11 this gives us something to work with that may be useful

terenzreignz (terenzreignz):

With permission amistre-sensei (^.^) If it covered 11 kilometers in one hour, then the time it takes trotting and walking together should add up to one hour, right? So consider x + y = 1 and make a system.

OpenStudy (amistre64):

9km/hr * n hrs = 9n km since we are looking for 11 km traveled, and what times were troted and sleeping ....

OpenStudy (amistre64):

of course this is just an idea i have :)

OpenStudy (amistre64):

thanks to euler x = 22/3 + 5n y = -11/3 - 3n would accomodate us and i just got your concern yes, x+y add up to an hour

OpenStudy (amistre64):

so our limits for n are such that 0 <= x,y <= 1

OpenStudy (amistre64):

22/3 + 5n > 0 ; n> -22/15 22/3 + 5n < 1 ; n< -19/15 y = -11/3 - 3n >0 ; n < -11/9 y = -11/3 - 3n <1 ; n > 2/9 the union of these 2 sets should define all the n parts that would work. with any luck :)

OpenStudy (amistre64):

y = -11/3 - 3n >0 ; n < -11/9 y = -11/3 - 3n <1 ; n > -20/9 looks better me thinks

terenzreignz (terenzreignz):

Why can't we just solve the system? T.T

OpenStudy (amistre64):

you can if you want too, but i really dont need the practice doing it that way ;)

terenzreignz (terenzreignz):

I see. <bows to sensei> (^.^)

terenzreignz (terenzreignz):

Yeah :P That was the gist of it :) We were looking for x (the number of hours the horse walked) and you got 2/3. So that's 2/3 hours which is exactly 40 minutes. Now as to how far the horse walked, the horse walked 9km/hr so just multiply that to 2/3, and you'll get the distance traveled. 2/3 of 9 is 6 (km) There ya go :P

OpenStudy (amistre64):

22/3 + 5n > 0 5n > -22/3 n > -22/15 22/3 + 5n < 1 5n < 1 - 22/3 n < 1/5 - 22/15 n < -19/15 -11/3 - 3n >0 - 3n > 11/3 n < -11/9 -11/3 - 3n <1 - 3n <1 + 11/3 n > -1/3 - 11/9 n > -14/9 lcm(9,15) = 45 -66/45 < n < -57/45 -70/45 < n < -55/45 looks like (-66/45, -57/45) is the interval for n to me

OpenStudy (amistre64):

x = 22/3 + 5(-66/45) x = 66/9 -66/9 = 0 y = -11/3 - 3(-66/45) y = -55/15 + 66/15 = 11/15 9(0) + 15(11/15) = 11 x = 22/3 + 5(-57/45) x = 66/9 -57/9 = 1 y = -11/3 - 3(-57/45) y = -55/15 + 57/15 = 2/15 9(1) + 15(2/15) = 11

OpenStudy (amistre64):

i know somethings amiss in it, but that was fun to try :)

terenzreignz (terenzreignz):

'Tis the realm of the gods :D

OpenStudy (amistre64):

hmmm, since x+y = 1 i could have done 22/3 + 5n -11/3 - 3n = 1 11/3 +2n = 1 2n = 1 - 11/3 n = -8/6 = -4/3 x = 22/3 +5(-4/3) = 22/3 - 20/3 = 2/3 , with any luck that means y = 1/3 y = -11/3 -3(-4/3) = -11/3 +12/3 = 1/3 :)